# The Value of Information in Human-AI Decision-making

### v1

@@ 3.2 Instance-level Agent-Complementary
Given a decision task with payoff function $S$ and an information model $\pi$, we define the instance-level agent-complementary information value of the realization $V=v$ as:
$ILIV^{\pi,S}(v;D^{b})=\mathrm{R}^{\pi_{v},S}(D^{b}\cup Z_{v})-\mathrm{R}^{\pi_{v},S}(D^{b})$
This local measure captures how much additional payoff can be gained by incorporating the specific realization $v$ into the agent decisions $D^{b}$. Summing $ILIV$ over all possible realizations of $V$ recovers the global agent-complementary information value ($ACIV$).
Proposition 3.4.
$ACIV^{\pi,S}(V;D^{b})=\sum_{v\in\mathbf{V}}ILIV^{\pi,S}(v;D^{b})$
We apply $ILIV$ to define an information-based explanation technique (ILIV-SHAP) that extends SHAP to explains how the information value of AI predictions complements human decisions $D^{\text{Human}}$ for specific instances. Vanilla SHAP (Lundberg and Lee, 2017) defines a saliency-based explanation with a set of effect variables $\phi_{i}$ representing the influence of the realization of basic signal $\Sigma_{i}=\sigma_{i}$ on the model output $f(v)$, where $v=(\sigma_{1},\ldots,\sigma_{n})$ and $f:\mathbf{V}\rightarrow\mathbb{R}$. Lundberg and Lee (2017) show that the only possible explanation model fulfills properties of local accuracy, missingness and consistency for model $f$ and instance $v$ is following:
Definition 3.5 (SHAP (Lundberg and Lee, 2017)).
$\phi_{i}(f,v)=\sum_{v^{\prime}\subseteq v}\frac{|v^{\prime}|!(n-|v^{\prime}|-1)!}{n!}[g_{f}(v^{\prime})-g_{f}(v^{\prime}\backslash\sigma_{i})]$
where $g_{f}(v^{\prime})$ denotes the expectation of model output conditioned on the subset of signals with $V^{\prime}=v^{\prime}$, i.e., $\text{\bf E}\![{f(V)|V^{\prime}=v^{\prime}}]$.
Instead of quantifying changes in the AI’s raw predictions, ILIV-SHAP clarifies how each feature contributes to the agent-complementary information value of predictions.
Proposition 3.6 (ILIV-SHAP).
An explanation model whose effective variable defined as
$\phi_{i}(ILIV(f(\cdot);D^{b}),v)$
fulfill properties of local accuracy, missingness and consistency for $ILIV$ of $f(\cdot)$ on instance $v$.
Intuitively, when $\phi_{i}(ILIV(f(\cdot);D^{b}),v)$ for a feature $\sigma_{i}$ is relatively large, it means that the model $f(\cdot)$ is extracting information that $D^{b}$ lacks from $\Sigma_{i}$ (or other features with equivalent information to it). Otherwise, when $\sigma_{i}$ makes a small relative contribution, it means that it makes the model ignore information in other features that have higher information value for human decision-makers (such that on average lossing information value).
@@ 3.3 Robustness Analysis of Information Order
Ambiguity about the appropriate payoff function is not uncommon in human-AI decision settings due to challenges of eliciting utility functions and potential variance in these functions across decision-makers or groups of instances; e.g., doctors penalize certain false negative results differently when diagnosing younger versus older patients (Mclaughlin and Spiess, 2023). We therefore define the partial order of complementary information value using Blackwell order. Blackwell’s comparison of signals (Blackwell et al., 1951) defines a (set of) signal $V_{1}$ as more informative than $V_{2}$ if $V_{1}$ has a higher information value on all possible decision problems. We identify this partial order by decomposing the space of decision problems via a basis of proper scoring rules  (Li et al., 2022; Kleinberg et al., 2023).
@@
A (set of) signal $V_{1}$ is Blackwell more informative than $V_{2}$ if $V_{1}$ achieves a higher payoff on any decision problem,
@@
We test the Blackwell order between signals on a basis of proper scoring rules induced from decision problems. The basis is the set of V-shaped scoring rules, parameterized by the kink of the piecewise-linear utility function.
@@
If
$\mathrm{R}^{\pi,S_{\mu}}(V_{1})\geq\mathrm{R}^{\pi,S_{\mu}}(V_{2}),\forall\mu\in(0,1),$
@@
$ACIV^{\pi,S_{\mu}}(V_{1};D^{b})\geq ACIV^{\pi,S_{\mu}}(V_{2};D^{b}),\forall\mu\in(0,1)$
This definition allows us to rank signals (or sets of signals) universally, without needing to pin down any specific payoff function prior to the analysis. See the use case in Section 4.
4 Experiment I: Model Comparison on Chest Radiographs Diagnosis
We apply our framework to a well-known cardiac dysfunction diagnosis task (Rajpurkar et al., 2018; Tang et al., 2020; Shreekumar, 2025). We apply the framework to analyze how much complementary information value a set of possible AI models offer to human decision-makers.
Figure 1: Information value of all deep-learning models calculated under our framework using Brier score as payoff function.
@@ 4.1 Data and Model
We use data from the MIMIC dataset (Goldberger et al., 2000), which contains anonymized electronic health records from Beth Israel Deaconess Medical Center (BIDMC), a large teaching hospital in Boston, Massachusetts, affiliated with Harvard Medical School. Specifically, we utilize chest x-ray images and radiology reports from the MIMIC-CXR database (Johnson et al., 2019) merged with patient and visit information from the broader MIMIC-IV database (Johnson et al., 2023). The payoff-related state, cardiac dysfunction $\omega\in\{0,1\}$, is coded based on two common tests, the NT-proBNP and the troponin, using the age-specific cutoffs from Mueller et al. (2019) and Heidenreich et al. (2022). We use the labels from Irvin et al. (2019) as the human decisions (without AI’s assistance) in the diagnosis task, which is a rule-based tool labeling the symptoms as positive, negative, or uncertain, i.e., $d\in\{+,?,-\}$ [^2] The three symbols represents the encoding we use for signal construction, not an assertion of how radiologists communicate.. We fune-tuned five deep-learning models on the cardiac dysfunction diagnosis task, VisionTransformer (Alexey, 2020), SwinTransformer (Liu et al., 2021), ResNet (He et al., 2016), Inception-v3 (Szegedy et al., 2016), and DenseNet (Huang et al., 2017). Our training set contains 12,228 images and validation set contains 6,115 images. On a hold-out test set with 12,229 images, the AUC achieved by the five models are: DenseNet with $0.77$, Inception v3 with $0.76$, ResNet with $0.77$, SwinTransformer with $0.78$, VisionTransformer with $0.80$.
We consider Brier score as a payoff function, and also conduct a robust analysis considering various V-shaped payoff functions with different kinks on a discretized grid of $[0,1]$ with a step $0.01$. We use the hold-out test set to estimate the data-generating process, which defines the joint distribution of state, human decisions and AI models’ predictions.
@@ 4.2 Results
Can the AI models complement human judgment? We first analyze the agent-complementary information values in Figure 1, using the Brier score as the payoff function. We find that all AI models provide complementary information value to human judgment. This highlights the same takeaways as section 5, that the AI model has considerable potential to improve human decisions. As shown in Figure 1 (comparison between $\mathrm{R}^{\pi,S}(D^{\text{Human}}\cup D^{\text{AI}})$ and $\mathrm{R}^{\pi,S}(D^{\text{Human}})$), all AI models capture at least $20\%$ of the total available information value that is not exploited by human decisions. This motivates deploying at AI to assist humans in this scenario.
From the other direction, the human decisions also provide complementary information to all AI models, comparing $\mathrm{R}^{\pi,S}(D^{\text{AI}}\cup D^{\text{AI}})$ with $\mathrm{R}^{\pi,S}(D^{\text{AI}})$ in Figure 1. In scenarios where partial automation is possible, this observation might inspire, for example, further investigation into what information the humans may have access to that is not represented in AI training data.
Which AI model offers the most decision-relevant information over human judgments? Figure 1 shows that VisionTransformer contains slightly higher information value than the other models and Inception v3 contains slightly lower information value than the other models. However, these differences are slight. If the payoff function were questionable in any way, an organization may not want to trust the model rankings. This motivates evaluating model performance over many possible losses to test if there is a Blackwell ordering of models. Across all the V-shaped payoff functions, we find that VisionTransformer is Blackwell more informative and Inception v3 is Blackwell less informative than all other models. By Proposition 3.9, we test the payoff of models on all V-shaped scoring rules, shown in Figure 4. The VisionTransformer achieves a higher information value on all V-shaped scoring rules, implying a higher information value on all decision problems. This analysis highlights the insufficiency of accuracy-based model comparisons to account for all downstream decision problems: 1) while accuracy may rank models in one way, there may exist decision problems where the order between models is reversed; 2) while two models may seem comparable in accuracy, one can be more informative than the other for all decision problems, which is robustly good.
@@ 5 Experiment II: Behavioral Analysis on Deepfake Detection
We analyze a deepfake video detection task (Dolhansky et al., 2020), where participants are asked to judge whether a video was created by generative AI. We apply the framework to benchmark the use of available information by the human, AI, and human-AI team.
Figure 2: Information value calculated under our framework in the information model defined by experiment data from Groh et al. (2022). The baseline on the left represents the expected payoff given no information, i.e., $\mathrm{R}^{\pi,S}(\emptyset)$, and the benchmark on the right represents the expected payoff given all available information, i.e., $\mathrm{R}^{\pi,S}(\{\Sigma_{1},\ldots,\Sigma_{n},D^{\text{Human}},D^{\text{Human-AI}},D^{\text{AI}}\})$. All the payoffs are calculated by $\mathrm{R}^{\pi,S}(\cdot)$, where $\cdot$ is the signal on the y-axis.
@@ 5.1 Data and Model
We define the information model on the experiment data of Groh et al. (2022). They recruited 5,524 non-expert participants through Prolific. Participants were asked to examine the videos, and provided with assistance from a computer vision model, which achieved $65\%$ accuracy on a holdout dataset. They reported their decisions in two rounds. They first reviewed the video and reported an initial decision ($D^{\text{Human}}$) without access to the model. Then, in a second round, they were told the AI’s recommendation ($D^{\text{AI}}$) and chose whether to change their initial decision, resulting in a final decision ($D^{\text{Human-AI}}$). Participants’ decisions (both initial and final) were elicited as a percentage indicating how confident they were that the video was a deepfake, measured in $1\%$ increments: $d\in\{0\%,1\%,\ldots,100\%\}$. We round the predictions from the AI to the same 100-scale probability scale available to study participants.
We use the Brier score as the payoff function: $S(\omega,d)=1-(\omega-d)^{2}$, with the binary payoff-related state: $\omega\in\{0,1\}=\{\text{genuine},\text{fake}\}$. The scale of Brier score is $[0,1]$ and a random guess ($d\sim\text{Bernoulli}(0.5)$) achieves $0.75$ payoff under Brier score. We choose the Brier score instead of the norm-1 accuaracy used by Groh et al. (2022) because Brier score is a proper scoring rule where truthfully reporting the belief maximizes the payoff [^3] We prefer a proper scoring rule so that the rational decision-maker’s strategy is to reveal their true belief, ensuring that the signal’s information value accurately reflects its role in forming beliefs. If the goal is merely to evaluate the outcome performance of real decision-makers, a non-proper scoring rule may still suffice..
We identify a set of features that were implicitly available to all three agents (human, AI, and human-AI). Because the video signal is high dimensional, we make use of seven seven video-level features using manually coded labels by Groh et al. (2022): graininess, blurriness, darkness, presence of a flickering face, presence of two people, presence of a floating distraction, and the presence of an individual with dark skin, all of which are labeled as binary indictors. These are the basic signals in our framework. We estimate the data-generating process $\pi$ using the realizations of signals, state, first round human decisions, AI predictions, and second round human-AI decisions.

### v3

@@ 3.2 Instance-level Agent-complementary Information Value
Given a decision task with payoff function $S$ and an information model $\pi$, we define the instance-level agent-complementary information value ($\operatorname{ILIV}$) of signal $V=v$ on instances where $V=v$ as:
$\operatorname{ILIV}^{v,\pi,S}(v;D^{b})=\mathrm{r}^{v,\pi,S}(v;D^{b})-\mathrm{r}^{v,\pi,S}(\emptyset;D^{b})$ (9)
where $\mathrm{r}^{v,\pi,S}(\emptyset;D^{b})$ represents the rational DM’s expected payoff on instances where $V=v$ with only the exisiting information encoded in agent decisions. Taking the expectation of $\operatorname{ILIV}$ over $V$ recovers the global agent-complementary information value ($\operatorname{ACIV}$), i.e.,
$\operatorname{ACIV}^{\pi,S}(V;D^{b})=\text{\bf E}_{v\sim\pi(v)}\![{\operatorname{ILIV}^{v,\pi,S}(v;D^{b})}]$
3.2.1 ILIV-SHAP Information-based Explanation
We apply $\operatorname{ILIV}$ to define an information-based explanation technique (ILIV-SHAP) that extends the canonical SHAP feature saliency explanation of a model’s prediction. ILIV-SHAP communicates how the data features lead to changes in the information value of AI predictions. While traditional SHAP summarizes the average contribution of each feature to a specific prediction over the baseline (prior) prediction, ILIV-SHAP summarizes the average contribution of each feature to the decision-relevant information value contained in the prediction.
Suppose a model $f$ that takes as input $m$ features and outputs a real number. Given an instance $\mathbf{x}=(x_{1},\ldots,x_{m})$, the importance of one feature $x_{i}$ to the model output $f(\mathbf{x})$ is encoded by the expected difference of model outputs when $x_{i}$ is marginalized out. Specifically, this is quantified by $f(\mathbf{x})-\text{\bf E}\![{f(X)|X_{-i}=\mathbf{x}_{-i}}]$, where $X_{-i}$ denotes all features except $X_{i}$. Considering the interaction between features, SHAP (Lundberg and Lee, 2017) uses the Shapley value to quantify the importance scores averaged on different combinations of features:
$\phi_{i}(f,\mathbf{x})=\sum_{\mathbf{x}^{\prime}\subseteq\mathbf{x}}\frac{|\mathbf{x}^{\prime}|!(m-|\mathbf{x}^{\prime}|-1)!}{m!}[g_{f}(\mathbf{x}^{\prime})-g_{f}(\mathbf{x}^{\prime}\backslash x_{i})]$
where $g_{f}(\mathbf{x}^{\prime})$ denotes the expected model output conditioned on $\mathbf{x}^{\prime}$, i.e., $\text{\bf E}\![{f(X)|X^{\prime}=\mathbf{x}^{\prime}}]$ for any $\mathbf{x}^{\prime}\subseteq\mathbf{x}$. The scores $\phi_{i}(f,\mathbf{x})$ output by SHAP construct an explanation model for a model output, which quantifies the expected counterfactual change in the model output caused by the feature $x_{i}$.
ILIV-SHAP extends SHAP to give an explanation model on how features impact the decision-relevant information value of an individual model output.
Definition 3.4 (ILIV-SHAP).
Given a model $f$ and data features $\mathbf{x}=(x_{1},\ldots,x_{m})$, the importance score of the i-th feature by ILIV-SHAP is
$\phi_{i}^{\operatorname{ILIV}}(f,\mathbf{x})=\sum_{\mathbf{x}^{\prime}\subseteq\mathbf{x}}\frac{|\mathbf{x}^{\prime}|!(m-|\mathbf{x}^{\prime}|-1)!}{m!}[\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime});D^{b})-\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime}\backslash x_{i});D^{b})]$
where $\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime});D^{b})$ denotes a counterfactual evaluation of $\operatorname{ILIV}$, which quantifies the expected payoff gain from additionally knowing $g_{f}(\mathbf{x}^{\prime})$ on the instances where $f(\mathbf{x})$ is realized. This counterfactual version of $\operatorname{ILIV}$ is guaranteed to achieve the maximum at $\mathbf{x}^{\prime}$ where $f(\mathbf{x})=g_{f}(\mathbf{x}^{\prime})$, i.e., the features missing from $\mathbf{x}^{\prime}$ have no impact on $f(\mathbf{x})$.
The explanation model offered by ILIV-SHAP is grounded in the following two properties. First, $\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$ are consistent with the extent to which feature $x_{i}$ contributes to the information value in $f(\mathbf{x})$. Specifically, for any two models $f$ and $f^{\prime}$ and any $\mathbf{x}^{\prime}\subseteq\mathbf{x}$, if $x_{i}$ contributes more to the information value in $f^{\prime}$ than the information value in $f$, i.e., $\operatorname{ILIV}^{f^{\prime}(\mathbf{x}),\pi,S}(g_{f^{\prime}}(\mathbf{x}^{\prime});D^{b})-\operatorname{ILIV}^{f^{\prime}(\mathbf{x}),\pi,S}(g_{f^{\prime}}(\mathbf{x}^{\prime}\backslash x_{i});D^{b})\geq\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime});D^{b})-\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(g_{f}(\mathbf{x}^{\prime}\backslash x_{i});D^{b})$, then $\phi^{\operatorname{ILIV}}_{i}(f^{\prime},\mathbf{x})\geq\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$. This property of ILIV-SHAP follows the consistency property of SHAP (Lundberg and Lee, 2017). Second, summing up the scores $\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$ recovers the information value of model output, $\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(f(\mathbf{x});D^{b})$. Formally, for any $f$ and any $\mathbf{x}$, $\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(f(\mathbf{x});D^{b})=\operatorname{ILIV}^{f(\mathbf{x}),\pi,S}(\text{\bf E}\![{f(X)}];D^{b})+\sum_{i=1}^{m}\phi^{\operatorname{ILIV}}_{i}(f,\mathbf{x})$. We demonstrate use of ILIV-SHAP in Section 6.
@@ 3.3 Robustness Analysis of Information Order
Our approach assumes the specification of a decision problem on which agents’ decisions and use of information are evaluated. However, ambiguity around the appropriate decision problem, and in particular, the appropriate scoring rule, is not uncommon in human-AI decision settings. Ambiguity can arise as a result of challenges in eliciting utility functions and/or variance in these functions across decision-makers or groups of instances; for example, doctors penalize certain false negative results differently when diagnosing younger versus older patients (Mclaughlin and Spiess, 2023). Blackwell’s comparison of signals (Blackwell et al., 1951) is an appropriate tool for addressing ambiguity about the payoff function, as it defines a (set of) signal $V_{1}$ as more informative than $V_{2}$ if $V_{1}$ has a higher information value on all possible decision problems. We identify this partial order by decomposing the space of decision problems via a basis of proper scoring rules [^2] For rational DMs, any decision problem can be represented by an equivalent proper scoring rule in Equation 1, such that the partial order defined via proper scoring rules also applies to the corresponding decisin tasks. (Li et al., 2022; Kleinberg et al., 2023).
@@
A signal $V_{1}$ is Blackwell more informative than $V_{2}$ if $V_{1}$ achieves a higher best-attainable payoff on any decision problems:
@@
The Blackwell order is evaluated over all possible decision problems, which cannot be tested directly. Fortunately, we only need to test over all proper scoring rules since any decision problem can be represented by an equivalent proper scoring rule, and the space of proper scoring rules can be characterized by a set of V-shaped scoring rules. A V-shaped scoring rule is parameterized by the kink of the piecewise-linear utility function.
@@
If $\forall\mu\in(0,1)$
$\mathrm{R}^{\pi,S_{\mu}}(V_{1})\geq\mathrm{R}^{\pi,S_{\mu}}(V_{2}),$
@@
$\operatorname{ACIV}^{\pi,S_{\mu}}(V_{1};D^{b})\geq\operatorname{ACIV}^{\pi,S_{\mu}}(V_{2};D^{b}),\forall\mu\in(0,1)$
This definition allows us to rank signals (or sets of signals) without needing to commit to a specific payoff function. We present a use case in Section 4.
4 Experiment I: Model Comparison on Chest Radiograph Diagnosis
We apply our framework to a well-known cardiac dysfunction diagnosis task (Rajpurkar et al., 2018; Tang et al., 2020; Shreekumar, 2025). We demonstrate how our framework can be used in model evaluation for analyzing how much complementary information value a set of possible AI models offers to the radiology reports written by experts.
Figure 1: Information value of all deep-learning models calculated under our framework. The first row represents the human-alone decisions (without considering any AI predictions as additional signals). The other rows are the combinations of the human-alone decisions and the AI predictions from different pre-trained models. We list the AI predictions alone to show the AI-complementary information value offered by human decisions.
@@ 4.1 Data and Model
We use data from the MIMIC dataset (Goldberger et al., 2000), which contains anonymized electronic health records from Beth Israel Deaconess Medical Center (BIDMC), a large teaching hospital in Boston, Massachusetts, affiliated with Harvard Medical School. Specifically, we utilize chest x-ray images and radiology reports from the MIMIC-CXR database (Johnson et al., 2019) merged with patient and visit information from the broader MIMIC-IV database (Johnson et al., 2023). The payoff-related state, cardiac dysfunction $\omega\in\{0,1\}$, is coded based on two common tests, the NT-proBNP and the troponin, using the age-specific cutoffs from Mueller et al. (2019) and Heidenreich et al. (2022). We label the radiology reports by a rule-based tool (Irvin et al., 2019) and use the labels as the human decisions (without AI assistance) in the diagnosis task to solve the problem of computational feasibility with high-dimensional textual reports. The labels are represented by the symptoms as positive, negative, or uncertain, i.e., $d\in\{+,?,-\}$. We fine-tuned five deep-learning models on the cardiac dysfunction diagnosis task, VisionTransformer (Alexey, 2020), SwinTransformer (Liu et al., 2021), ResNet (He et al., 2016), Inception-v3 (Szegedy et al., 2016), and DenseNet (Huang et al., 2017). Our training set contains 12,228 images, and the validation set contains 6,115 images. On a hold-out test set with 12,229 images, the AUC achieved by the five models is: DenseNet with $0.77$, Inception v3 with $0.76$, ResNet with $0.77$, SwinTransformer with $0.78$, and VisionTransformer with $0.80$.
We consider Brier score, a.k.a., quadratic score, as the payoff function: $S(\omega,d)=1-(\omega-d)^{2}$. The scale of the quadratic score is $[0,1]$ and a random guess ($d\sim\text{Bernoulli}(0.5)$) achieves $0.75$ payoff. We use the quadratic score instead of the mean absolute error that is usually used in cardiac dysfunction diagnosis task because the quadratic score is a proper scoring rule where truthfully reporting the belief maximizes the payoff [^3] We prefer a proper scoring rule so that the rational decision-maker’s strategy is to reveal their true belief, ensuring that the signal’s information value accurately reflects its role in forming beliefs.. We also conduct a robust analysis considering various V-shaped payoff functions with different kinks on a discretized grid of $[0,1]$ with a step of $0.01$. We use the hold-out test set to estimate the data-generating process, which defines the joint distribution of state, human decisions, and AI models’ predictions.
We construct the scale of performances by a no-information bound and a full-information bound. The no-information bound is $\mathrm{R}^{\pi,S}(\emptyset)$, the baseline as we define the information value. The full-information bound is defined as the expected payoff of a rational DM who has access to all signals, human label from radiology report and predictions from five AI models. We do not consider the radiology images as a basic signal in this task because of the problem of computational feasibility with high-dimensional signals. We use the five vision models’ predictions as embeddings of the original images. See more detail on the algorithm of choosing the optimal embeddings in Appendix B.
@@ 4.2 Results
Can the AI models complement human judgment?
We first analyze the agent-complementary information values in Figure 1, using the Brier score as the payoff function. We find that all AI models provide complementary information value to the aforementioned human judgment. As shown in Figure 1 (comparison between $\mathrm{R}^{\pi,S}(D^{\text{Human}}\cup D^{\text{AI}})$ and $\mathrm{R}^{\pi,S}(D^{\text{Human}})$), all AI models capture at least $20\%$ of the total available information value (across all AI model and human decisions) that is not exploited by human decisions. This motivates deploying an AI to assist humans in this scenario.
In the other direction, the human decisions also provide complementary information to all AI models, comparing $\mathrm{R}^{\pi,S}(D^{\text{Human}})$ with $\mathrm{R}^{\pi,S}(D^{\text{AI}})$ in Figure 1. This observation might inspire, for example, further investigation of the information the humans can access to that is not represented in AI training data.
Which AI model offers the most decision-relevant information over human judgments?
Figure 1 shows that VisionTransformer contains slightly higher information value than the other models, and Inception v3 contains slightly lower information value than the other models. We assess the stability of VisionTransformer’s superiority over the other AI models across many possible losses to test if there is a Blackwell ordering of models. By Proposition 3.7, we test the payoff of models on all V-shaped scoring rules, shown in Figure 4. Across all the V-shaped payoff functions, we find that VisionTransformer is Blackwell more informative and Inception v3 is Blackwell less informative than all other models. The VisionTransformer achieves a higher information value on all V-shaped scoring rules, implying a higher information value on all decision problems.
@@ 5 Experiment II: Behavioral Analysis on Deepfake Detection
We apply our framework to analyze a deepfake video detection task (Dolhansky et al., 2020), where participants are asked to judge whether a video was created by generative AI, including with the assistance of an AI model.
Figure 2: Information value calculated under our framework in the information model defined by the experiment of Groh et al. (2022). Basic signals include the seven video level features and three types of agent decisions. The baseline on the left represents the expected payoff given no information, i.e., $\mathrm{R}^{\pi,S}(\emptyset)$, and the benchmark on the right represents the expected payoff given all available information, i.e., $\mathrm{R}^{\pi,S}(\{\Sigma_{1},\ldots,\Sigma_{n},D^{\text{Human}},D^{\text{Human-AI}},D^{\text{AI}}\})$. All the payoffs are calculated by $\mathrm{R}^{\pi,S}(\cdot)$, where $\cdot$ is the signal on the y-axis.
@@ 5.1 Data and Model
We define the information model on the experiment data of Groh et al. (2022). Non-expert participants (n=5,524) were recruited through Prolific and asked to examine the videos. They reported their decisions in two rounds. They first reviewed the video and reported an initial decision ($D^{\text{Human}}$) without access to the AI model. Then, in a second round, they were provided with the recommendation ($D^{\text{AI}}$) of a multitask cascaded convolutional neural network (Zhang et al., 2016), with estimated $65\%$ accuracy on holdout data, and chose whether to change their initial decision. This produced a final decision ($D^{\text{Human-AI}}$). Both decisions were elicited as a percentage indicating how confident the participant was that the video was a deepfake, measured in $1\%$ increments: $d\in\{0\%,1\%,\ldots,100\%\}$. We round the predictions from the AI model to the same 100-scale probability scale available to study participants.
We use the Brier score as the payoff function, with the binary payoff-related state: $\omega\in\{0,1\}=\{\text{genuine},\text{fake}\}$. This choice differs from the mean absolute error used by Groh et al. (2022), but again we use the quadratic score because it is a proper scoring rule where truthfully reporting the belief maximizes the score.
We identify a set of features that were implicitly available to all three agents (human, AI, and human-AI). Because the video signal is high dimensional, we make use of seven video-level features using manually coded labels by Groh et al. (2022): graininess, blurriness, darkness, presence of a flickering face, presence of two people, presence of a floating distraction, and the presence of an individual with dark skin, all of which are labeled as binary indictors. These are the basic signals in our framework. We estimate the data-generating process $\pi$ using the realizations of signals, state, first-round human decisions, AI predictions, and second-round human-AI decisions. The no-information bound is the same as Section 4 while the full-information bound is defined as the expected payoff of a rational DM who has access to all video-level features and three agents’ decisions.